An argon-ion laser emits a blue-green beam of light with a wavelength of 488nm in a vacuum.
Compute the energy in joules for the two energy states for the atomic transition which produces
this light?
1
Expert's answer
2020-06-09T13:21:05-0400

Energy is calculated by "\frac{hc}{\lambda}=\frac{6.63\times10^{-34}\times3\times10^8}{488\times10^{-9}}=4\times10^{-19}joules"